### v1

@@ 5.3 The role of CO in the $2175\,\AA$ bump
In the dust accretion rate we introduced a limit to the growth of carbonaceous grains argued on the basis of rapid formation of CO molecules at high gas densities (for $n>10^{3}\,\rm H\,cm^{-3}$, see section 3.4). To explore the impact of CO, we have also tested a range of models where we neglect the role of the removal of C atoms into CO molecules: i.e. both carbonaceous and silicate grains stop their growth by accretion of refractory material at densities $n>10^{4}\,\rm H\,cm^{-3}$ due to icy mantles coating the grain’s surface.
@@
The effect on the depletion of C and on the extinction curve is shown in Fig. 21, resp. top and bottom panels, for several galaxy masses and for different gas fractions. Removing the consideration on CO formation for the subgrid model of carbonaceous grain growth results in a larger depletion of C, i.e. more carbonaceous grains are grown from the gas phase. Consequently, the $2175\,\AA$ bump strength of the extinction curve is more prominent as a result of a large amount of small carbonaceous grains due to their stronger growth by accretion, which favors the growth of small grains. In addition, without considering CO formation, the extinction curves are also shallower in their UV-to-optical slope due to the larger ratio of carbonaceous to silicate grains, since there is, proportionally, a lower amount of small silicate grains that are responsible for the steepness of the slope. A striking behavior with CO modeling and its impact on carbonaceous dust growth is that galaxies with the low gas fractions exhibit weak variation of the amount of (small) carbonaceous grains with respect to the CO modeling, as opposed to simulated galaxies with a high fraction of gas. Indeed galaxies with high gas fractions have much more CO formed compared to galaxies with low gas fractions as they reach high gas surface densities, are more turbulent and, hence, get more gas that can reach typical threshold of CO formation at $10^{3}\,\rm H\,cm^{-3}$. We can also see that for the simulations with the largest difference in bump strength due to the absence of CO corresponds also a significant shift in the UV-to-optical slope: there is overall more carbonaceous grains with respect to silicate grains, which produces a shallower slope of the extinction curve.
@@
Figure 22 illustrates this last point by showing the estimated CO gas surface densities $\Sigma_{\rm CO}$ for two simulated galaxies of the same mass and with the same initial gas metallicity but which differ in initial gas fraction (G9LG_VLZ versus G9HG_VLZ). Due its larger gas fraction, the G9HG_VLZ run produces much more CO than the G9LG_VLZ run, with CO concentrated in regions of high gas densities. G9HG_VLZ contains $\simeq 60$ times more CO than G9LG_VLZ, while their total gas mass ratio are only $\simeq 3$. Since gas accretion primarily proceeds in dense gas, the effect of C capture in CO molecular formation is a strong limiter to growth of carbonaceous grains, and is, hence, the stronger in galaxies with large gas densities. The estimated CO masses are naive as we only assume instantaneous processing of the entire C gas mass into molecules at a threshold density: it does not take into account the time it takes to grow the CO or destroy it, nor the effect of destruction by UV radiation field and by cosmic rays. We defer such investigation and its impact on carbonaceous grain growth to future studies.
@@ 6.1 Polycyclic aromatic hydrocarbons
The presence of multiple emission bands in the mid-IR is well attested and the vibrational excitation of small organic molecules, made of PAHs, is the favorite picture to explain these emission features (Leger & Puget 1984; Desert et al. 1990; Draine & Li 2007; Tielens 2013). Observations of the 3.3/11.3 $\mu$m band ratio near photo-dissociation regions suggests that, within the Milky Way ISM, PAHs are primarily composed from molecules with $\sim 50-100$ C atoms (Croiset et al. 2016), well below the centroid of our small grain bin. Despite their small size, PAHs can lock up to $\sim 10$% of C in the ISM. They are hypothesised to play a key role in the thermo-chemistry of the ISM, by heating HI gas and the warm phase via the photo-ejection of electrons (e.g. Bakes & Tielens 1994; Weingartner & Draine 2001b), catalyzing the production of molecular hydrogen, or acting as the formation sites of small hydrocarbons (see the review by Kaiser et al. 2015, and references therein).
Theoretical modelling of PAH survival (subject to the diffuse ISM radiation field, SN shocks and thermal sputtering) implies that efficient mechanisms for their replenishment are necessary to account for their abundance (Micelotta 2009), may this be either through a ‘top-bottom’ picture via the shattering of small and/or large carbonaceous grains (Seok et al. 2014; Hirashita & Murga 2020; Narayanan et al. 2023), or from combustion-like reaction in the ISM giving rise to the most fundamental aromatic species (e.g. benzene, naphtalene, Reizer et al. 2022). Observations imply a significantly more linear relation of PAH mass with galaxy metallicity (e.g. Madden et al. 2006; Galliano et al. 2021; Shivaei et al. 2024), and a strong correlation of PAH band emission with the interstellar radiation field (e.g. Pilleri et al. 2012; Egorov et al. 2023), supporting that there are ISM processes that alter differently the PAH and the dust grain abundances. Of fundamental relevance for the present work, PAHs are also candidate sources of the $2175\,\AA$ feature. However, in our extinction curve modelling, the origin of this feature comes from the graphitic properties assumed for our small carbonaceous grains ($5\,\rm nm$) with the Weingartner & Draine (2001a) optical constants. Therefore, a model in which this feature is originated instead by smaller, less resilient particles (e.g. PAHs), could result in a larger variation (compared to the one presented here) of the $2175\,\AA$ bump strength with ISM conditions and galaxy metallicity.
@@ 6.2 Dust in stellar ejecta
We have seen that for low metallicity galaxies ($\gtrsim 0.1\,\rm Z_{\odot}$), their extinction curves are reflective of their inability to grow carbonaceous dust by gas accretion, hence, their dust grain composition is skewed towards silicates, and carbonaceous dust is only composed of big grains shaped by the SNII ejecta. For extremely low metallicities galaxies ($\lesssim 0.1\,\rm Z_{\odot}$), the gas accretion onto silicate grains is also severely quenched, and SNII ejecta become the only source of both carbonaceous and silicate dust growth in the ISM. Dust condensation efficiencies for SNe are highly uncertain (Schneider & Maiolino 2023). Dust in SNe condenses in the shell of the young ejecta, with dust masses ranging $[0.03-1]\,\rm M_{\odot}$ for $20\,\rm M_{\odot}$ progenitor (e.g. Sarangi & Cherchneff 2015; Marassi et al. 2019; Brooker et al. 2022), before the destruction of dust takes on by the ignition of the SN reverse shock through sputtering and grain-grain collisions (see e.g. Kirchschlager et al. 2019), which settles the final amount of dust that is going to be mixed with the ISM. How much of the dust survives in this final phase depends on the magnetic field, initial dust size distribution, and clumpiness of the ejecta. Therefore, final dust masses remain largely uncertain, with dust survival rates ranging from a few % to a few 10 %, and with large uncertainties in the final dust size distributions (e.g. Silvia et al. 2010; Micelotta et al. 2016; Kirchschlager et al. 2019, 2023; Slavin et al. 2020). On the observational side, dust mass estimates are also affected by strong uncertainties related to dust composition, size distribution and temperature, with best estimates for Cassiopeia A and Crab Nebula that are in the range $[0.1-1]\,\rm M_{\odot}$ and $[0.025-0.05]\,\rm M_{\odot}$, respectively (see the compilation of data in and the discussion in Schneider & Maiolino 2023). Given all those uncertainties in dust yields, it is difficult to predict accurately the dust composition, and, hence, the extinction curves from low mass and low metallicity galaxies. As we have seen in our simulations, the contribution from ejecta is of critical importance in $Z\gtrsim 0.1\,\rm Z_{\odot}$, as opposed that of silicate grains in that same metallicity range. To gauge the impact of the uncertainties on dust ejecta, we have run some of the G8HG low mass galaxies with different fraction of small grain size fraction in stellar ejecta $f_{\rm ej,S}$ in appendix B. Values of the extinction curve UV-to-optical slope and bump strength are significantly increased with higher $f_{\rm ej,S}$ as long as the grain growth of a given species is dominated by the contribution from stellar ejecta. If carbonaceous grains are the direct product of SN ejecta in low mass galaxies as suggested by these simulations, the extinction curves of the Magellanic Clouds, and young galaxies of the early Universe (Witstok et al. 2023; Markov et al. 2024) through their attenuation curves, could serve as indirect constraints on mean values of condensation efficiencies in SNII ejecta.
@@ 6.3 Approximations in the grain properties
An important foundation of the model at the heart of the obtained grain size distribution, especially for the MRN-like size distribution in the Milky Way analogue, is the velocity dispersion of grains. We have assumed, as in other work (e.g. Aoyama et al. 2017; Gjergo et al. 2018; Granato et al. 2021), a unique mean grain velocity dispersion for coagulation and a unique $\sigma_{\rm gr}$-$n$ scaling relation for shattering, although different in magnitude for the small and big grains (for respectively coagulation and shattering rates). However, it is clear that the grain velocity dispersion should vary with the grain properties, such as their charge, and with the ISM properties, its density, ionisation, turbulence, and magnetization (Yan et al. 2004; Moseley et al. 2023). Additionally, the velocity dispersion of grains should imply a grain-gas decoupling for the largest grain sizes at the molecular cloud scale, with a typical diffusion length of $L\simeq 5\,a_{0.1}(n/1\,{\rm cm^{-3}})^{-1}(\sigma_{\rm gr}/c_{\rm s})\,\rm pc$, with the consequence that global shattering rates should increase. We defer such investigations with, using e.g. a diffusion-like approach (Lebreuilly et al. 2019), to future work.
We have adopted a coarse two-size grain decomposition, as opposed to decompose fully the size spectrum of grains based on the model of Hirashita (2015) (a similar approach that has also been employed in Hou et al. 2017; Aoyama et al. 2017; Gjergo et al. 2018; Granato et al. 2021). This approach significantly reduces the memory and computational requirements of the model, as it only necessitates 4 bins (2 sizes, 2 chemical compositions) instead of the $2N_{\rm bin}$, where $N_{\rm bin}\gtrsim 10$ used in practice for full size spectrum decomposition (McKinnon et al. 2018; Aoyama et al. 2019; Li et al. 2021). Furthermore, the two-size bin approximation greatly simplifies the treatment of coagulation and shattering that otherwise would have needed cross-coupling between size bins, which makes the computations of coagulation and shattering rates to scale like $N_{\rm bin}^{2}$. While a detailed description of fine-sampled grain size distribution associated with accurate models of grain size dispersion could be considered a gold standard, this might come at the expense of spatial resolution, which is crucial for capturing the multiphase ISM responsible for phase transitions between accretion/coagulation-dominated regions (in dense neutral or molecular gas) and shattering-dominated regions (in diffusion ionized gas).
@@ 7 Conclusion
We have introduced a novel model of dust growth for galaxy simulations in the ramses code that is coupled to the chemical enrichment of the gas. This model decomposes dust into carbonaceous and silicate grains with two grain sizes of $5\,\rm nm$ and $0.1\,\rm\mu m$. It includes the modeling of dust growth by accretion, stellar ejecta, dust destruction by SNe, thermal sputtering and astration, and also includes coagulation and sputtering to allow dust to cycle between small and large grain sizes. Using hydrodynamical simulations of idealised disc galaxies with a multiphase ISM structure captured at a resolution of $20\,\rm pc$, we have explored various galaxy masses, metallicities, gas fractions, and their effect on the galactic dust properties and extinction curves.
@@
Those simulations have been performed in a idealized context without considering the effect of fresh gas replenishment by halo-scale gas accretion. A natural direct extension of this work is to test the validity of the dust model in fully cosmological hydrodynamical simulations to understand the influence of cosmic infall and mergers on the properties of dust. In particular, such cosmological simulations could be used to understand the cosmological build up of dust in galaxies, and test the fraction of galaxies with obscured star formation (Magnelli et al. 2020; Pozzi et al. 2021; Fudamoto et al. 2021; Inami et al. 2022). In this work, we have limited ourselves to testing extinction curves of simulated galaxies, which gives a direct insight on dust properties. Although attenuation curves are the result of the complex interplay between the dust properties, and the ISM and stellar distributions (Narayanan et al. 2018), the flurry of observational evidence might serve as an important testbed for the cosmic evolution of dust properties.
@@ Appendix C Sputtering times
Figure 25 compares the sputtering times as a function of the gas temperature for a grain size of $a=0.1\,\mu\rm m$ and a gas density of $n=1\,\rm H\,cm^{-3}$ from Tielens et al. (1994), Tsai & Mathews (1995), and  Hu et al. (2019). Hu et al. (2019) provides a convenient fifth-order polynomial fit to the predictions of the sputtering rates $Y(T)$ from Nozawa et al. (2006). The values of the Hu et al. (2019) fit with sufficient digits are reported in table 3 (private communication with C.-Y. Hu). The formula from Tsai & Mathews (1995) is intermediate to the sputtering times of carbonaceous and silicate grains from Tielens et al. (1994) and Hu et al. (2019), which both show a general good agreement up to a temperature of $T\sim 10^{8}\,\rm K$. Note that these calculations are all for a semi-infinite target, however, the finite size of the grain modifies the sputtering rates for grains with size smaller than the penetration depth of the impinging ion (e.g. Kirchschlager et al. 2019).

### v2

@@ 5.3 The role of CO in the $2175\,\AA$ bump
In the dust accretion rate we introduced a limit to the growth of carbonaceous grains argued on the basis of rapid formation of CO molecules at high gas densities (for $n>10^{3}\,\rm H\,cm^{-3}$, see section 3.4). To explore the impact of CO, we also tested a range of models where we neglect the role of the removal of C atoms into CO molecules: i.e. both carbonaceous and silicate grains stoped their growth by accretion of refractory material at densities $n>10^{4}\,\rm H\,cm^{-3}$ due to icy mantles coating the grain’s surface.
@@
The effect on the depletion of C and on the extinction curve is shown in Fig. 21, resp. top and bottom panels, for several galaxy masses and for different gas fractions. Removing the consideration on CO formation for the subgrid model of carbonaceous grain growth results in a larger depletion of C, i.e. more carbonaceous grains are grown from the gas phase. Consequently, the $2175\,\AA$ bump strength of the extinction curve is more prominent as a result of a large amount of small carbonaceous grains due to their stronger growth by accretion, which favours the growth of small grains. In addition, without considering CO formation, the extinction curves are also shallower in their UV-to-optical slope due to the larger ratio of carbonaceous to silicate grains, since there is, proportionally, a lower amount of small silicate grains that are responsible for the steepness of the slope. A striking behaviour with CO modelling and its impact on carbonaceous dust growth is that galaxies with the low gas fractions exhibit weak variation of the amount of (small) carbonaceous grains with respect to the CO modelling, as opposed to simulated galaxies with a high fraction of gas. Indeed galaxies with high gas fractions have much more CO formed compared to galaxies with low gas fractions as they reach high gas surface densities, are more turbulent and, hence, get more gas that can reach typical threshold of CO formation at $10^{3}\,\rm H\,cm^{-3}$. We can also see that for the simulations with the largest difference in bump strength due to the absence of CO corresponds also a significant shift in the UV-to-optical slope: there is overall more carbonaceous grains with respect to silicate grains, which produces a shallower slope of the extinction curve.
@@
Figure 22 illustrates this last point by showing the estimated CO gas surface densities $\Sigma_{\rm CO}$ for two simulated galaxies of the same mass and with the same initial gas metallicity but which differ in initial gas fraction (G9LG_VLZ versus G9HG_VLZ). Due its larger gas fraction, the G9HG_VLZ run produces much more CO than the G9LG_VLZ run, with CO concentrated in regions of high gas densities. G9HG_VLZ contains $\simeq 60$ times more CO than G9LG_VLZ, while the ratio of their total gas mass is only $\simeq 3$. Since gas accretion primarily proceeds in dense gas, the effect of C capture in CO molecular formation is a strong limiter to growth of carbonaceous grains, and is, hence, the stronger in galaxies with large gas densities. The estimated CO masses are naive as we only assume instantaneous processing of the entire C gas mass into molecules at a threshold density: it does not take into account the time it takes to grow the CO or destroy it, nor the effect of destruction by UV radiation field and by cosmic rays. We defer such investigation and its impact on carbonaceous grain growth to future studies.
@@ 6.1 Polycyclic aromatic hydrocarbons
The presence of multiple emission bands in the mid-IR is well attested and the vibrational excitation of small organic molecules, made of PAHs, is the favourite picture to explain these emission features (Leger & Puget 1984; Desert et al. 1990; Draine & Li 2007; Tielens 2013). Observations of the 3.3/11.3 $\mu$m band ratio near photo-dissociation regions suggests that, within the Milky Way ISM, PAHs are primarily composed from molecules with $\sim 50-100$ C atoms (Croiset et al. 2016), well below the centroid of our small grain bin. Despite their small size, PAHs can lock up to $\sim 10$% of C in the ISM. They are hypothesised to play a key role in the thermo-chemistry of the ISM, by heating HI gas and the warm phase via the photo-ejection of electrons (e.g. Bakes & Tielens 1994; Weingartner & Draine 2001b), catalysing the production of molecular hydrogen, or acting as the formation sites of small hydrocarbons (see the review by Kaiser et al. 2015, and references therein).
Theoretical modelling of PAH survival (subject to the diffuse ISM radiation field, SN shocks and thermal sputtering) implies that efficient mechanisms for their replenishment are necessary to account for their abundance (Micelotta 2009), may this be either through a ‘top-bottom’ picture via the shattering of small and/or large carbonaceous grains (Seok et al. 2014; Hirashita & Murga 2020; Narayanan et al. 2023), or from combustion-like reaction in the ISM giving rise to the most fundamental aromatic species (e.g. benzene, naphthalene, Reizer et al. 2022). Observations imply a significantly more linear relation of PAH mass with galaxy metallicity (e.g. Madden et al. 2006; Galliano et al. 2021; Shivaei et al. 2024), and a strong correlation of PAH band emission with the interstellar radiation field (e.g. Pilleri et al. 2012; Egorov et al. 2023), supporting that there are ISM processes that alter differently the PAH and the dust grain abundances. Of fundamental relevance for the present work, PAHs are also candidate sources of the $2175\,\AA$ feature. However, in our extinction curve modelling, the origin of this feature comes from the graphitic properties assumed for our small carbonaceous grains ($5\,\rm nm$) with the Weingartner & Draine (2001a) optical constants. Therefore, a model in which this feature is originated instead by smaller, less resilient particles (e.g. PAHs), could result in a larger variation (compared to the one presented here) of the $2175\,\AA$ bump strength with ISM conditions and galaxy metallicity.
@@ 6.2 Dust in stellar ejecta
We have seen that for low metallicity galaxies ($\gtrsim 0.1\,\rm Z_{\odot}$), their extinction curves are reflective of their inability to grow carbonaceous dust by gas accretion, hence, their dust grain composition is skewed towards silicates, and carbonaceous dust is only composed of big grains shaped by the SNII ejecta. For extremely low metallicities galaxies ($\lesssim 0.1\,\rm Z_{\odot}$), the gas accretion onto silicate grains is also severely quenched, and SNII ejecta become the only source of both carbonaceous and silicate dust growth in the ISM. Dust condensation efficiencies for SNe are highly uncertain (Schneider & Maiolino 2023). Dust in SNe condenses in the shell of the young ejecta, with dust masses ranging $[0.03-1]\,\rm M_{\odot}$ for $20\,\rm M_{\odot}$ progenitor (e.g. Sarangi & Cherchneff 2015; Marassi et al. 2019; Brooker et al. 2022), before the destruction of dust takes on by the ignition of the SN reverse shock through sputtering and grain-grain collisions (see e.g. Kirchschlager et al. 2019), which settles the final amount of dust that is going to be mixed with the ISM. How much of the dust survives in this final phase depends on the magnetic field, initial dust size distribution, and clumpiness of the ejecta. Therefore, final dust masses remain largely uncertain, with dust survival rates ranging from a few % to a few 10 %, and with large uncertainties in the final dust size distributions (e.g. Silvia et al. 2010; Micelotta et al. 2016; Kirchschlager et al. 2019, 2023; Slavin et al. 2020). On the observational side, dust mass estimates are also affected by strong uncertainties related to dust composition, size distribution and temperature, with best estimates for Cassiopeia A and Crab Nebula that are in the range $[0.1-1]\,\rm M_{\odot}$ and $[0.025-0.05]\,\rm M_{\odot}$, respectively (see the compilation of data and the discussion in Schneider & Maiolino 2023).
Given all those uncertainties in dust yields, it is difficult to predict accurately the dust composition and, hence, the extinction curves from low mass and low metallicity galaxies. As we have seen in our simulations, the contribution from ejecta is of critical importance in $Z\gtrsim 0.1\,\rm Z_{\odot}$, as opposed that of silicate grains in that same metallicity range. To gauge the impact of the uncertainties on dust ejecta, we ran some of the G8HG low mass galaxies with different fraction of small grain size fraction in stellar ejecta $f_{\rm ej,S}$ in appendix B. Values of the extinction curve UV-to-optical slope and bump strength are significantly increased with higher $f_{\rm ej,S}$ as long as the grain growth of a given species is dominated by the contribution from stellar ejecta. If carbonaceous grains are the direct product of SN ejecta in low mass galaxies as suggested by these simulations, the extinction curves of the Magellanic Clouds, and young galaxies of the early Universe (Witstok et al. 2023; Markov et al. 2024) through their attenuation curves, could serve as indirect constraints on mean values of condensation efficiencies in SNII ejecta.
@@ 6.3 Approximations in the grain properties
An important foundation of the model at the heart of the obtained grain size distribution, especially for the MRN-like size distribution in the Milky Way analogue, is the velocity dispersion of grains. We assumed, as in other work (e.g. Aoyama et al. 2017; Gjergo et al. 2018; Granato et al. 2021), a unique mean grain velocity dispersion for coagulation and a unique $\sigma_{\rm gr}$-$n$ scaling relation for shattering, although different in magnitude for the small and big grains (for respectively coagulation and shattering rates). However, it is clear that the grain velocity dispersion should vary with the grain properties, such as its charge, and with the ISM properties, such as its density, ionisation, turbulence, and magnetisation (Yan et al. 2004; Moseley et al. 2023). Additionally, the velocity dispersion of grains should imply a grain-gas decoupling for the largest grain sizes at the molecular cloud scale, with a typical diffusion length of $L\simeq 5\,a_{0.1}(n/1\,{\rm cm^{-3}})^{-1}(\sigma_{\rm gr}/c_{\rm s})\,\rm pc$, with the consequence that global shattering rates should increase. We defer such investigations with, using e.g. a diffusion-like approach (Lebreuilly et al. 2019), to future work.
We adopted a coarse two-size grain decomposition, as opposed to decompose fully the size spectrum of grains based on the model of Hirashita (2015) (a similar approach that is employed in Hou et al. 2017; Aoyama et al. 2017; Gjergo et al. 2018; Granato et al. 2021). This approach significantly reduces the memory and computational requirements of the model, as it only necessitates 4 bins (2 sizes, 2 chemical compositions) instead of the $2N_{\rm bin}$, where $N_{\rm bin}\gtrsim 10$ used in practice for full size spectrum decomposition (McKinnon et al. 2018; Aoyama et al. 2019; Li et al. 2021). Furthermore, the two-size bin approximation greatly simplifies the treatment of coagulation and shattering that otherwise would have needed cross-coupling between size bins, which makes the computations of coagulation and shattering rates to scale like $N_{\rm bin}^{2}$. While a detailed description of fine-sampled grain size distribution associated with accurate models of grain size dispersion could be considered a gold standard, this might come at the expense of spatial resolution, which is crucial for capturing the multiphase ISM responsible for phase transitions between accretion/coagulation-dominated regions (in dense neutral or molecular gas) and shattering-dominated regions (in diffuse ionised gas).
@@ 7 Conclusion
We introduced a novel model of dust growth for galaxy simulations in the ramses code that was coupled to the chemical enrichment of the gas. This model decomposed dust into carbonaceous and silicate grains with two grain sizes of $5\,\rm nm$ and $0.1\,\rm\mu m$. It included the modelling of dust growth by accretion, stellar ejecta, dust destruction by SNe, thermal sputtering and astration, and also included coagulation and sputtering to allow dust to cycle between small and large grain sizes. Using hydrodynamical simulations of idealised disc galaxies with a multiphase ISM structure captured at a resolution of $20\,\rm pc$, we explored various galaxy masses, metallicities, gas fractions, and their effect on the galactic dust properties and extinction curves.
@@
Those simulations were carried out in an idealised context without considering the effect of fresh gas replenishment by halo-scale gas accretion. A natural direct extension of this work is to test the validity of the dust model in fully cosmological hydrodynamical simulations to understand the influence of cosmic infall and mergers on the properties of dust. In particular, such cosmological simulations could be used to understand the cosmological build up of dust in galaxies, and test the fraction of galaxies with obscured star formation (Magnelli et al. 2020; Pozzi et al. 2021; Fudamoto et al. 2021; Inami et al. 2022). In this work, we limited ourselves to testing extinction curves of simulated galaxies, which gives a direct insight on dust properties. Although attenuation curves are the result of the complex interplay between the dust properties, and the ISM and stellar distributions (Narayanan et al. 2018), the flurry of observational evidence might serve as an important testbed for the cosmic evolution of dust properties.
@@ Appendix C Sputtering times
Figure 25 compares the sputtering times as a function of the gas temperature for a grain size of $a=0.1\,\mu\rm m$ and a gas density of $n=1\,\rm H\,cm^{-3}$ from Tielens et al. (1994), Tsai & Mathews (1995), and  Hu et al. (2019). Hu et al. (2019) provides a convenient fifth-order polynomial fit to the predictions of the sputtering rates $Y(T)$ from Nozawa et al. (2006). The values of the Hu et al. (2019) fit with sufficient digits are reported in table 3 (private communication with C.-Y. Hu). The formula from Tsai & Mathews (1995) is intermediate to the sputtering times of carbonaceous and silicate grains from Tielens et al. (1994) and Hu et al. (2019), which both show a general good agreement up to a temperature of $T\sim 10^{8}\,\rm K$. We note that these calculations are all for a semi-infinite target, however, the finite size of the grain modifies the sputtering rates for grains with size smaller than the penetration depth of the impinging ion (e.g. Kirchschlager et al. 2019).